# Continuous-time quantum walk-based ansätze on neutral atom hardware

### v1

@@ II.5.2 Bracelet states
where $\tau_{\rm{min}}^{\rm{hw}}$ is the minimum allowed walk time ($\sim 0.4$) and the 2-2 guards against quantization error arising through hardware constraints so the walk time are fixed and only the phasor unitary $\vec{\gamma}=\{\gamma_{k}\}_{k=1}^{p}$ remain variational. As the resulting landscape is non-convex, we optimize these with the derivative-free COBYLA optimizer initialized at $\gamma_{k}=0$, stopping the scan over peaks when there is a decrease in success probability or the range of identified peaks is exhausted.
We discard the first peak because the state is concentrated in the $C_{N}$-symmetric (translation-invariant) subspace. Since the phasor unitary depends on Hamming weight, it acts as $e^{i\gamma h}\mathbb{I}$ on each weight-$h$ sector and therefore cannot distinguish distinct $D_{N}$-symmetric components. The total walk time $\tau_{\rm{tot}}$ must be long enough to accumulate relative phase between $D_{N}$-symmetric components at fixed $h$ before optimization over $\vec{\gamma}$ can appreciably improve the success probability.
More generally, the required $\tau_{\rm eff}$ for high-fidelity preparation of $|[z^{*}]\rangle$ is explainable with a frequency-resolution model. Let $\{(\lambda_{r},|r\rangle)\}_{r=1}^{d}$ be the eigenpairs of the walk generator $\mathcal{G}$ restricted to the $D_{N}$-invariant subspace [36]. Because the only interleaved control is the phasor unitary, resolvability of the target state is governed by spectral gaps $\lambda_{rs}=|\lambda_{r}-\lambda_{s}|$ between eigenmodes $r,s\in\{1,\dots,d\}$ that both have appreciable support on the same Hamming-weight sector as the target [37]. Consequently, control under variation of $\vec{\gamma}$ requires order-unity accumulated phase, $\lambda_{rs}\,\tau_{\mathrm{eff}}=\Theta(1)$.
@@
The bottom panel of Fig. 3 shows that, across all distinct targets $|[z^{*}]\rangle$ at $N=5$ to $12$, the accumulated walk time $\tau_{\rm{eff}}$ scales linearly with $1/\Delta_{\rm min}(\kappa^{*})$ with a single fixed $\kappa^{*}$. Within each $N$, targets that are a maximal independent set consistently exhibit the smallest $\tau_{\rm eff}$ and the largest $\Delta_{\min}$. This aligns with prior work on scarred quantum walks [38], which places these states on or near a small, low-degree hypercube-like subgraph weakly coupled to the rest of the graph – with dynamics confined to this subgraph producing sparser, quasi-regular spectra with larger internal gaps, resulting in faster preparation.
Figure 3: Bracelet state preparation. Top: Population of the $N=9$ target bracelet state $|[z^{\rm half}]\rangle$ during a continuous-time quantum walk from $|0\rangle$, sampled at $\tau_{j}=j\,\Delta\tau$ with $\Delta\tau=0.02$, over $\tau\in[0,20]$. Local maxima $\tau^{(m)}_{\rm{peak}}$ above the threshold $1/|D_{N}|$ are marked by a pink ‘X’, which serve as fixed parameters in variational optimization of phases $\gamma$. Bottom: Accumulated walk time $\tau_{\mathrm{eff}}$ required to reach success probability $\geq 0.98$ against the inverse resolvable spectral minimum $1/\Delta_{\rm{min}}(\kappa^{*})$ (see Eqs. 36 and 37) for all distinct targets $|[z^{*}]\rangle$ with $N=5,\dots,12$. Here $\Delta_{\rm{min}}(\kappa)$. We fix $\kappa^{*}\approx 7.2\pm 0.4$ from a stability plateau by maximizing $\mathrm{mean}(r)-2\,\mathrm{std}(r)$ over sliding $\kappa$ windows and taking the window center. Maximum-independent-set (MIS) targets are circled in black.
@@ III Rydberg atom arrays
{python}
import numpy as np import bloqade.analog as ba # QuEra’s SDK
# — Input parameters — N = 20 # Set the number of atoms tau = 2*np.pi*1.2345 # Unitless evolution time Omega = 2*np.pi*2.5 # Aquila’s Rabi frequency # —
# Evolution time in usec evolution_time = tau / Omega # Dynamic blockade radius in um d = (5_420_503/Omega)**(1/6) D = d/(2*np.sqrt( np.sin(np.pi/N)*np.sin(2*np.pi/N)))
D *= 1.0 # Optionally add a variational # fudge factor to optimize the blockade radius.
theta = np.linspace(0,2*np.pi,N+1)[0:N] positions = np.array([D*np.sin(theta), D*np.cos(theta)]).T positions = [tuple(q) for q in positions] positions = ba.atom_arrangement.
ListOfLocations(positions) rabi_drive = ba.constant( duration=evolution_time, value=Omega) detuning = ba.constant( duration=evolution_time, value=0.0) program = ba.rydberg_h( atoms_positions=positions, amplitude=rabi_drive, detuning=detuning) # Run using emulation data = program.bloqade.python().run(100) # Run using Aquila via BraKet data2 = program.braket.aquila().run_async(100)
@@ IV Results: Preparation of product and bracelet states
We present results for the preparation of the product states $z^{\rm half}$ and $z^{\rm MIS}$, together with their corresponding bracelet states $z^{\rm half}]$ and $z^{\rm MIS}]$, as described in Section II.3, using the two phase-walk ansätze introduced in Section II.5. These are realized in three scenarios: the “perfect” CTQW dynamics, its approximation by the Rydberg Hamiltonian in noiseless emulation, and experimental implementation on Aquila, sharing the same variational parameters across each case. Our comparison focuses on the success probability and amplification relative to a uniform baseline, contrasting the observed amplification as a function of both the cardinality of the invariant (blockaded) subspace $|\mathcal{V}|$ and the ansatz depth $p$, following the power-law scaling relationships introduced in Section II.4. Particular attention is given to the extent to which the scaling behavior predicted by the ideal CTQW is qualitatively reproduced on noisy hardware.
@@
Figure 5: Probability distributions for preparation of the $z^{\rm half}$ product state (top panel) and the corresponding bracelet state (middle panel) at $N=7$, obtained from numerical simulation of the ideal CTQW dynamics (“Perfect”), noiseless emulation of the Rydberg Hamiltonian (“Emulation”), the raw probability distribution from 1000 shots on Aquila, and the reconstructed distribution (“EM”) that accounts for measurement error. The table below summarizes the corresponding success probabilities: the State column labels the target state; Ansatz Depth is the number of walk layers; the Naïve count column gives the fraction of bitstrings matching the target directly from raw counts; the Perfect column reports the preparation probability under ideal CTQW dynamics; the Emulation column shows the same quantity using noiseless Rydberg-atom emulation. The EM (H) column gives the success probability from Aquila hardware data after mitigation of measurement errors, while EM (E) reports the corresponding value from 1000 emulated shots convolved with the noisy measurement channel. Data for other target states are shown in tables 2 and 3.
@@
Figure 6: Comparison of success probabilities $P(z^{*})$ (top row) and amplification factors $A=|\mathcal{V}|\,P(z^{*})$ (bottom row) for $z^{\rm half}$ (left panels) and $z^{\rm MIS}$ (right panels) across CTQW (Perfect), noiseless Rydberg emulation, and Aquila hardware. Results are shown at depths $p=1,2,3$ for $z^{\rm half}$ and $p=1,2$ for $z^{\rm MIS}$. Amplification plots include power-law fits to Eq. 13 (dashed lines). For $z^{\rm half}$ targets, CTQW fits yield $c=\{1.099,\,1.112,\,1.015\}$ with exponents $\alpha=\{0.878\!\pm\!0.006,\,0.935\!\pm\!0.003,\,0.981\!\pm\!0.002\}$ ($R^{2}=\{0.9870,\,0.9997,\,0.9997\}$); emulation gives $c=\{0.632,\,0.719,\,0.785\}$ with $\alpha=\{0.877\!\pm\!0.009,\,0.934\!\pm\!0.006,\,0.930\!\pm\!0.006\}$ ($R^{2}=\{0.9881,\,0.9961,\,0.9959\}$); and hardware yields $c=\{0.595,\,0.463,\,0.418\}$ with $\alpha=\{0.729\!\pm\!0.041,\,0.876\!\pm\!0.025,\,0.734\!\pm\!0.051\}$ ($R^{2}=\{0.9326,\,0.9374,\,0.7889\}$). For $z^{\rm MIS}$ targets, CTQW remains near-linear with $c=\{0.995,\,1.000\}$ and $\alpha=\{0.996\!\pm\!0.001,\,1.000\!\pm\!0.000\}$ ($R^{2}=\{0.9972,\,1.0000\}$); emulation gives $c=\{0.641,\,0.507\}$ with $\alpha=\{1.000\!\pm\!0.004,\,0.967\!\pm\!0.006\}$ ($R^{2}=\{0.9838,\,0.9857\}$); and hardware yields $c=\{0.512,\,0.285\}$ with $\alpha=\{0.820\!\pm\!0.029,\,0.703\!\pm\!0.067\}$ ($R^{2}=\{0.9724,\,0.9250\}$).
@@
Figure 5 presents the unprocessed probability distributions for the $z^{\rm half}$ product and bracelet states at $N=7$ across the three considered scenarios, together with the success probabilities obtained from $1000$ shots in noiseless emulation convolved with the asymmetric bit-flip channel, and from the same number of shots on Aquila. Here, and for all other considered target states (see tables , 2 and 3), the reconstructed distributions obtained from the scrambled emulation results agree with the true values within the 95% confidence intervals.
@@ IV.1 Product state preparation
[…]
@@
Overall, these results are consistent with the super-linear amplification scaling in $|\mathcal{V}|$ expected of an efficient phase-walk ansatz, and to a lesser extent with increasing depth. The phasor unitary relies on local detuning, which carries a shot-to-shot per-site coherent error of roughly 10%, leading to deviations from the intended $\pm\pi$ phase shifts that compound across iterations. This is likely the primary reason for the breakdown of increased convergence at higher $p$, in addition to the effects of incoherent errors from longer program times. It is therefore notable that the $z^{\rm half}$ targets still show improved convergence at $p=2$. For the $z^{\rm MIS}$ targets, the CTQW dynamics predict convergence near unity across the range of $N$. Consequently, the decrease in amplification from $p=1$ to $p=2$ in both emulated and experimentally implemented Rydberg dynamics is consistent with saturation of CTQW-based convergence at $p=1$, where further iterations offer little additional amplification and are outweighed by the loss in success probability from accumulated phase errors.
@@ IV.2 Bracelet state preparation
In bracelet state preparation, we find that direct parameter transfer from the ideal CTQW-based ansatz to its approximation via the Rydberg Hamiltonian is unreliable due to the effect of van der Waals interactions over $\tau_{\rm{eff}}$, which is significantly longer than the $\tau$ used in product state preparation. These interactions introduce an additional positive phase through Rydberg-Rydberg couplings, which contributes coherent phase error to the walk evolution and breaks the $\gamma\equiv\gamma$ equivalence that holds for the ideal CTQW. However, we find that this effect can be mitigated through joint optimization of $\tau_{\rm{eff}}$ and $\bm{\gamma}$. By taking as the objective the mean success probability
@@
we are able to identify parameter sets that achieve high convergence simultaneously in both the CTQW and Rydberg cases, with the global phasor (see Eq. 33) encoded into a piecewise constant phase profile of the Rabi drive.
@@
Compared to the CTQW-derived parameters, the jointly optimal parameters systematically require a longer walk time. In Fig. 8, the ratio of the jointly-optimal cumulative walk time $\tau_{\rm{eff}}^{\rm{Ryd}}$ over the CTQW optimal $\tau_{\rm eff}^{\rm CTQW}$ against the CTQW-predicted $1/\Delta_{\min}$ instances reveals that target states with smaller spectral gaps have the largest inflation in $\tau_{\rm eff}$ (up to $\sim 10\times$), while more “resolvable” targets cluster around a factor of $\sim 1.5$-$2.5$. This behavior indicates that van-der-Waals-induced phases compress the effective gap, thus requiring a longer effective walk time to resolve the target. The $\left[z^{\rm MIS}\right]$ states are more strongly impacted than the $\left[z^{\rm half}\right]$ targets, which is consistent with their higher Rydberg density and the correspondingly faster accumulation of interaction-induced phases.
@@
Figure 8: Comparison between the ideal CTQW-based bracelet-state preparation ansatz and its implementation via the Rydberg Hamiltonian for states $\left[z^{\rm half}\right]$ and $\left[z^{\rm MIS}\right]$ (circled in black) that achieved $\frac{1}{2}\left(P_{\rm CTQW}([z^{*}])+P_{\rm Ryd}([z^{*}])\right)\geq 0.8$ under joint parameter optimization. The horizontal axis shows the CTQW-predicted inverse resolvable spectral minimum $1/\Delta_{\min}$ (also shown in Fig. 3), while the vertical axis shows the ratio of optimal walk times found through the procedure described in Section II.5.2.
Figure 9: Performance for bracelet state preparation of $\left[z^{\rm half}\right]$ (top) and $\left[z^{\rm MIS}\right]$ (bottom) on Aquila (filled markers outlined in black, error bars denote 95% CI), compared with noiseless emulation (hollow markers) and ideal CTQW dynamics (half-filled markers). The left panels show success probability as a function of cumulative walk time $\tau_{\rm eff}$, while the right panels show the amplification of the target state as a function of $|\mathcal{V}|$.
In Fig. 9, we see that success probability for preparation of $\left[z^{\rm half}\right]$ and $\left[z^{\rm MIS}\right]$ on hardware decreases approximately linearly with $\tau_{\rm{eff}}$, which most strongly impact the $\left[z^{\rm half}\right]$ targets. Nevertheless, the amplification of $\left[z^{\rm MIS}\right]$ is qualitatively consistent with power-law scaling in $|\mathcal{V}|$, with the even $N$ instances positively offset from the odd, which may be explained by the even case corresponding to a superposition over two states. In contrast, the odd cases span an $N$-fold superposition. In contrast, while emulated results for the $\left[z^{\rm half}\right]$ are also consistent with a power-law scaling, the hardware results do not reproduce this trend, due to the accumulation of error associated with the longer required cumulative walk time. Although, since circuit depth scales linearly with $\tau_{\rm eff}$ in this setting, unlike in product-state preparation, we cannot preclude the existence of alternative parameter sets that achieve equivalent or superior performance at shorter effective walk times.

### v2

@@ II.6.2 Bracelet states
where $\tau_{\rm{min}}^{\rm{hw}}$ is the minimum allowed walk time ($\sim 0.4$) and the 2-2 guards against quantization error arising through hardware constraints so the walk time are fixed and only the phase separator $\vec{\gamma}=\{\gamma_{k}\}_{k=1}^{p}$ remain variational. As the resulting landscape is non-convex, we optimize these with the derivative-free COBYLA optimizer initialized at $\gamma_{k}=0$, stopping the scan over peaks when there is a decrease in success probability or the range of identified peaks is exhausted.
We discard the first peak because the state is concentrated in the $C_{N}$-symmetric (translation-invariant) subspace. Since the phase separator depends on Hamming weight, it acts as $e^{i\gamma h}\mathbb{I}$ on each weight-$h$ sector and therefore cannot distinguish distinct $D_{N}$-symmetric components. The total walk time $\tau_{\rm{tot}}$ must be long enough to accumulate relative phase between $D_{N}$-symmetric components at fixed $h$ before optimization over $\vec{\gamma}$ can appreciably improve the success probability.
More generally, the required $\tau_{\rm eff}$ for high-fidelity preparation of $|[z^{*}]\rangle$ is explainable with a frequency-resolution model. Let $\{(\lambda_{r},|r\rangle)\}_{r=1}^{d}$ be the eigenpairs of the walk generator $\mathcal{G}$ restricted to the $D_{N}$-invariant subspace [29]. Because the only interleaved control is the phase separator, resolvability of the target state is governed by spectral gaps $\lambda_{rs}=|\lambda_{r}-\lambda_{s}|$ between eigenmodes $r,s\in\{1,\dots,d\}$ that both have appreciable support on the same Hamming-weight sector as the target [44]. Consequently, control under variation of $\vec{\gamma}$ requires order-unity accumulated phase, $\lambda_{rs}\,\tau_{\mathrm{eff}}=\Theta(1)$.
@@
The bottom panel of Fig. 3 shows the accumulated walk time $\tau_{\rm{eff}}$ required to reach high-fidelity preparation against $1/\Delta_{\rm min}(\kappa^{*})$ for all dihedrally distinct bracelet targets $|[z^{*}]\rangle$ at $N=5$ to $12$. These numerical results are consistent with the scaling relationship predicted by the frequency-resolution model. Across all targets, $\tau_{\rm{eff}}$ scales linearly with $1/\Delta_{\rm min}(\kappa^{*})$ for a single fixed $\kappa^{*}$, consistent with Eq. 37. This $1/\Delta_{\rm min}$ scaling stands in contrast to adiabatic state preparation, where the required evolution time scales as $1/\Delta_{\rm min}^{2}$ [45]. The improvement arises because the phase-walk ansatz operates in a non-adiabatic regime, creating a superposition of eigenmodes whose relative phases evolve and interfere constructively on the target rather than adiabatically tracking a single instantaneous ground eigenstate.
Within each $N$, targets that are a maximal independent set consistently exhibit the smallest $\tau_{\rm eff}$ and the largest $\Delta_{\min}$. This is consistent with prior work on scarred quantum walks [46], which places these states on or near a small, low-degree hypercube-like subgraph weakly coupled to the rest of the graph – with dynamics confined to this subgraph producing sparser, quasi-regular spectra with larger internal gaps, resulting in faster preparation.
Figure 3: Bracelet state preparation. Top: Population of the $N=9$ target bracelet state $|[z^{\rm half}]\rangle$ during a continuous-time quantum walk from $|0\rangle$, sampled at $\tau_{j}=j\,\Delta\tau$ with $\Delta\tau=0.02$, over $\tau\in[0,20]$. Local maxima $\tau^{(m)}_{\rm{peak}}$ above the threshold $1/|D_{N}|$ are marked by a pink ‘X’, which serve as fixed parameters in variational optimization of phases $\gamma$. Bottom: Accumulated walk time $\tau_{\mathrm{eff}}$ required to reach success probability $\geq 0.98$ against the inverse resolvable spectral minimum $1/\Delta_{\rm{min}}(\kappa^{*})$ (see Eqs. 36 and 37) for all dihedrally distinct targets $|[z^{*}]\rangle$ with $N=5,\dots,12$. Here $\Delta_{\rm{min}}(\kappa)$. We fix $\kappa^{*}\approx 7.2\pm 0.4$ from a stability plateau by maximizing $\mathrm{mean}(r)-2\,\mathrm{std}(r)$ over sliding $\kappa$ windows and taking the window center. Maximum-independent-set (MIS) targets are circled in black.
@@ III Rydberg atom arrays
import bloqade.analog as ba # QuEra’s SDK
# --- Input parameters ---
N = 20 # Set the number of atoms
tau = 2*np.pi*1.2345 # Unitless evolution time
Omega = 2*np.pi*2.5 # Aquila’s Rabi frequency
# Evolution time in usec
evolution_time = tau / Omega
# Dynamic blockade radius in um
d = (5_420_503/Omega)**(1/6)
D = d/(2*np.sqrt(
np.sin(np.pi/N)*np.sin(2*np.pi/N)))
D *= 1.0 # Optionally add a variational
# fudge factor to optimize the blockade radius.
theta = np.linspace(0,2*np.pi,N+1)[0:N]
positions = np.array([D*np.sin(theta),
D*np.cos(theta)]).T
positions = [tuple(q) for q in positions]
positions = ba.atom_arrangement.\\
ListOfLocations(positions)
rabi_drive = ba.constant(
duration=evolution_time,
value=Omega)
detuning = ba.constant(
duration=evolution_time,
value=0.0)
program = ba.rydberg_h(
atoms_positions=positions,
amplitude=rabi_drive,
detuning=detuning)
# Run using emulation
data = program.bloqade.python().run(100)
# Run using Aquila via BraKet
data2 = program.braket.aquila().run_async(100)
@@ IV Results: Preparation of product and bracelet states
We present results for the preparation of the product states $z^{\rm half}$ and $z^{\rm MIS}$, together with their corresponding bracelet states $z^{\rm half}]$ and $z^{\rm MIS}]$, as described in Section II.4, using the two phase-walk ansätze introduced in Section II.6. These are realized in three scenarios: the “perfect” CTQW dynamics, its approximation by the Rydberg Hamiltonian in noiseless emulation, and experimental implementation on Aquila, sharing the same variational parameters across each case. Our comparison focuses on the success probability and amplification relative to a uniform baseline, contrasting the observed amplification as a function of both the cardinality of the invariant (blockaded) subspace $|\mathcal{V}|$ and the ansatz depth $p$, following the power-law scaling relationships introduced in Section II.5. Particular attention is given to the extent to which the scaling behavior predicted by the ideal CTQW is qualitatively reproduced on noisy hardware.
@@
Figure 5: Probability distributions for preparation of the $z^{\rm half}$ product state and the corresponding bracelet state at $N=7$, obtained from numerical simulation of the ideal CTQW dynamics (“Perfect”), noiseless emulation of the Rydberg Hamiltonian (“Emulation”), the raw probability distribution from 1000 shots on Aquila, and the reconstructed distribution (“EM”) that accounts for measurement error. Top shows the product state distribution. Middle shows the bracelet state distribution. The table below summarizes the corresponding success probabilities: the State column labels the target state; Ansatz Depth is the number of walk layers; the Naïve count column gives the fraction of bitstrings matching the target directly from raw counts; the Perfect column reports the preparation probability under ideal CTQW dynamics; the Emulation column shows the same quantity using noiseless Rydberg-atom emulation. The EM (H) column gives the success probability from Aquila hardware data after mitigation of measurement errors, while EM (E) reports the corresponding value from 1000 emulated shots convolved with the noisy measurement channel. Agreement between EM (E) and the Emulation column validates that the EM procedure correctly recovers the underlying distribution. Data for all target states are shown in Tables 2 and 3 of Appendix C.
@@
Figure 6: Comparison of success probabilities and amplification factors for $z^{\rm half}$ and $z^{\rm MIS}$ across CTQW (Perfect), noiseless Rydberg emulation, and Aquila hardware. Top shows success probabilities $P(z^{*})$, with $z^{\rm half}$ on the left and $z^{\rm MIS}$ on the right. Bottom shows amplification factors $A=|\mathcal{V}|\,P(z^{*})$. Results are shown at depths $p=1,2,3$ for $z^{\rm half}$ and $p=1,2$ for $z^{\rm MIS}$. Amplification plots include power-law fits to Eq. 13 (dashed lines); fit parameters are reported in Table 4.
@@
Figure 5 presents the unprocessed probability distributions for the $z^{\rm half}$ product and bracelet states at $N=7$ across the three considered scenarios, together with the success probabilities obtained from $1000$ shots in noiseless emulation convolved with the asymmetric bit-flip channel, and from the same number of shots on Aquila. Here, and for all other considered target states (see Tables 2 and 3), the reconstructed distributions obtained from the scrambled emulation results agree with the true values within the 95% confidence intervals.
@@ IV.1 Product state preparation
The walk parameters for the target $z^{\rm half}$ and $z^{\rm MIS}$ product states are obtained through ideal simulation of a quantum walk generator, that is, a walk without Rydberg interactions, blockade violations, and finite rise times. They are then directly transferred, with no further optimization, to the time-dependent Rydberg Hamiltonian as described above. The local phase separator (see Eq. 16) is encoded into a minimum-duration triangle pulse with a 5050\,ns rise and fall time commensurate with the constraints of Aquila [25], and a peak value of $\Delta_{\text{max}}=\phi/(50\,\text{ns})\leq 62\,\text{rad}/\mu\text{s}$.
@@
Overall, these results are consistent with the super-linear amplification scaling in $|\mathcal{V}|$ expected of an efficient phase-walk ansatz, and to a lesser extent with increasing depth. The phase separator relies on local detuning, which carries a shot-to-shot per-site coherent error of roughly 10%, leading to deviations from the intended $\pm\pi$ phase shifts that compound across iterations. This is likely the primary reason for the breakdown of increased convergence at higher $p$, in addition to the effects of incoherent errors from longer program times. It is therefore notable that the $z^{\rm half}$ targets still show improved convergence at $p=2$. For the $z^{\rm MIS}$ targets, the CTQW dynamics predict convergence near unity across the range of $N$. Consequently, the decrease in amplification from $p=1$ to $p=2$ in both emulated and experimentally implemented Rydberg dynamics is consistent with saturation of CTQW-based convergence at $p=1$, where further iterations offer little additional amplification and are outweighed by the loss in success probability from accumulated phase errors.
@@ IV.2 Bracelet state preparation
In bracelet state preparation, direct parameter transfer from the ideal CTQW-based ansatz to its approximation via the Rydberg Hamiltonian is unreliable due to the effect of van der Waals interactions over $\tau_{\rm{eff}}$, which is significantly longer than the $\tau$ used in product state preparation. These interactions introduce an additional positive phase through Rydberg-Rydberg couplings, contributing coherent phase error to the walk evolution and breaking the $\gamma\equiv\gamma$ equivalence that holds for the ideal CTQW. We mitigate this effect through joint optimization of $\tau_{\rm{eff}}$ and $\bm{\gamma}$, taking as the objective the mean success probability
@@
to identify parameter sets that achieve high convergence simultaneously in both cases, with the global phase separator (see Eq. 33) encoded into a piecewise constant phase profile of the Rabi drive.
@@
Compared to the CTQW-derived parameters, the jointly optimal parameters systematically require longer walk times. In Fig. 8, the ratio $\tau_{\rm{eff}}^{\rm{Ryd}}/\tau_{\rm eff}^{\rm CTQW}$ reveals that target states with smaller spectral gaps require the largest inflation in $\tau_{\rm eff}$, up to approximately $10\times$. More resolvable targets cluster around a factor of $1.5$ to $2.5$. This behavior indicates that van-der-Waals-induced phases compress the effective gap, requiring longer walk times to resolve the target.
@@
Figure 8: Comparison between the ideal CTQW-based bracelet-state preparation ansatz and its implementation via the Rydberg Hamiltonian for states $\left[z^{\rm half}\right]$ and $\left[z^{\rm MIS}\right]$ (circled in black) that achieved $\frac{1}{2}\left(P_{\rm CTQW}([z^{*}])+P_{\rm Ryd}([z^{*}])\right)\geq 0.8$ under joint parameter optimization. The horizontal axis shows the CTQW-predicted inverse resolvable spectral minimum $1/\Delta_{\min}$ (also shown in Fig. 3), while the vertical axis shows the ratio of optimal walk times found through the procedure described in Section II.6.2.
Figure 9 presents the amplification and fidelity scaling for bracelet state preparation, comparing ideal CTQW dynamics, noiseless Rydberg emulation, and Aquila hardware. Unlike the product state experiments, which sweep ansatz depth at fixed $\tau_{\rm eff}\approx\pi/2$, bracelet state preparation is performed at a single optimized $(\tau_{\rm eff},p)$ pair for each target, with $\tau_{\rm eff}$ ranging from $\sim\!5$ to $\sim\!20$ across targets (see Table 3 of Appendix C). The top panels show amplification (Eq. 12) as a function of graph size $|\mathcal{V}|$, with data split by the parity of $N$. For $\left[z^{\rm MIS}\right]$ (top right), power-law fits (Eq. 13) to the hardware data yield $\alpha=0.84\pm 0.07$ for even-$N$ instances and $\alpha=0.69\pm 0.08$ for odd-$N$, both indicating super-quadratic ($\alpha>0.5$) scaling. The even-$N$ instances correspond to a two-fold degenerate orbit ($|[z]|=2$) and exhibit near-linear scaling across all three cases, while odd-$N$ instances span an $N$-fold superposition and show somewhat reduced but still clear power-law scaling. For $\left[z^{\rm half}\right]$ (top left), noiseless emulation follows the CTQW prediction, but hardware results show no clear power-law trend ($\alpha=0.18\pm 0.18$), most likely due to the combined effect of longer walk times and greater ansatz depth.
[…]
The fitted $T_{2}^{*}$ is notably shorter than the single-qubit Rabi coherence time of $\sim\!5\,\mu$[25]. This reduction likely reflects the accumulation of phase errors across the many Rabi phase jumps. Because the ansatz depth scales with the required walk time (see Eq. 35), longer $\tau_{\rm eff}$ entails proportionally more phase jumps (p=10p=104848 across the targets studied here). Future work could explore whether fidelity on hardware might be improved by using fewer, longer walk segments.
Despite these limitations, the hardware results are consistent with the non-adiabatic $\tau_{\rm eff}\propto 1/\Delta_{\rm min}$ scaling predicted by the frequency-resolution model (Eq. 37). As shown in the bottom-right panel of Fig. 9, the optimized walk times cluster around the $\tau\propto 1/\Delta_{\rm min}$ reference line rather than the $\tau\propto 1/\Delta_{\rm min}^{2}$ scaling that would arise from adiabatic preparation [45]. While the van-der-Waals-induced phase compression increases the proportionality constant relative to the ideal CTQW (Fig. 8), the scaling relationship is preserved, confirming that the phase-walk ansatz successfully implements the non-adiabatic “shortcut” on current hardware.